OpenStudy (anonymous):

4(-3x+1)=-10(x-4)-14x ? help !

OpenStudy (anonymous):

12x + 4 = -10x + 40 - 14x equating common terms, 12x + 10x + 14x = 40 - 4 36x = 36 x = 1

OpenStudy (anonymous):

X=3

OpenStudy (anonymous):

3

OpenStudy (anonymous):

-12x+4=-10x+40-14x -12x+4=-24x+40 Then subtract 4 from 40 to equal 36 and add 24to -12 to equal 12. Divide both by 12 and x gives you 3.

OpenStudy (anonymous):

4

OpenStudy (anonymous):

THANK YOU GUYS ! :)